# Hybrid RAG-empowered Multi-modal LLM for Secure Healthcare Data Management: A Diffusion-based Contract Theory Approach

### v1

@@ VI Numerical Results
In this section, we conduct extensive experiments to assess the performance of the proposed hybrid RAG-empowered MLLM framework in particular healthcare analysis and evaluate the effectiveness of our proposed incentive mechanism. For MLLM inference, we use Python 3.10.14 running on CPU intel Xeon(R) Gold 6133 and GPU NVIDIA RTX A6000 to execute tasks. For the simulation setting of the diffusion model, we list the main parameter settings in III.
@@ VI-A Case Study of Hybrid RAG-empowered Medical MLLMs
We simulate a prototype of the hybrid RAG-empowered MMLMs with the support of LLaVA-Med[30] and llamaindex [^1] https://docs.llamaindex.ai/. As illustrated in Fig. 2, we present two examples to demonstrate the functionality and application of our framework. Upon receiving a task query with multi-modal healthcare data, the framework first retrieves the corresponding data from the respective modal database and performs a preliminary screening of $P$ results based on the cosine similarity between vectors. Next, hybrid RAG further refines these results using the MIS metric to identify the best matches, which are then used as inputs to the MLLMs. Finally, the MLLMs process all this information to provide diagnostic outputs and personalized services according to the query.
We apply the criteria of Responsive Artificial Intelligence (RAI) [^2] https://www.microsoft.com/en-us/ai/responsible-ai to assess whether the outputs of MLLMs present potential risks related to morality, bias, and ethics. Additionally, we also assess the relationship between the output of MLLMs and task query with the Semantic Similarity (SS)[35], which reflects the quality of the diagnosis and serves as a crucial indicator for measuring the MLLMs’ output. Due to the lack of evaluation benchmarks, we integrate LLM evaluators[36] with prompt engineering techniques [26] to measure the quality of data analysis for MLLMs under the method of GPT4-o, LLaVA, and LLaVA with Hybrid RAG. The scoring is normalized to a range of $[0,1]$. Higher scores denote greater reliability and lack of bias in the MLLM output, along with a strong correlation to the task query information. In contrast, lower scores signify a substantial gap from the anticipated results. Finally, we combine the RAI evaluation and SS into a unified score, known as the relative LLM score $\zeta$[37], which is calculated using the formula:
$\zeta=\lambda\cdot RAI+\nu\cdot SS,$ (20)
where $\lambda$ and $\nu$ are the weighting factors for the RAI and SS, respectively. In our approach, we assign equal weights by setting $\lambda=0.5$ and $\nu=0.5$.
As shown in Fig. 3, we present the performance comparison between the proposed framework under different healthcare data cases. Our findings indicate that hybrid RAG enables LLaVA-Med to consistently score above $0.9$, particularly in X-ray cases from Users 1 and 2 with known etiologies, maintaining high-quality answers and stability. In contrast, other MLLMs exhibit reduced output quality due to the interference of disease factors. In the scenarios involving Users 3 and 4, who are normal without specific causes, the MLLMs achieve high scores and deliver reasonable judgments. However, in the case of User 5, who is normal but has an X-ray that can easily be misjudged by a doctor, other MLLMs exhibit a higher misjudgment rate. In contrast, hybrid RAG continues to produce high-quality outputs by matching similar disease conditions. These cases illustrate that the retrieved data retrieved by hybrid RAG provides valuable information for answering questions. We summarize all the scores in Table IV, which clearly demonstrates that hybrid RAG helps LLAVA-Med maintain consistently high scores, showcasing its strong performance across different scenarios. This indicates that hybrid RAG effectively considers the quality of retrieved information by utilizing the features of multi-modal data, including images and texts. The retrieved relevant healthcare data can aid MLLMs through contextual relationships, allowing MLLMs to deliver reliable and robust outputs owing to their powerful contextual learning capabilities.
@@ VI-B Performance of GDM-based Contract Theory Approach
For the proposed contract model, we employ an on-policy GDM algorithm within a double actor-critic framework, and the specific settings of training hyperparameters are shown in Table III. In our setup, we consider $10$ healthcare data holders divided into two types, with $M=10$ and $K=2$. For the two types of healthcare data holders $\theta_{1}$ and $\theta_{2}$, values are randomly sampled from the intervals $(1,6)$ and $(13,18)$, respectively. Additionally, the maximum tolerance of AoI $\overline{A}_{max}$ is sampled randomly within the range of $(30,60)$. For the utility of the MLLM service provider, the parameters $\alpha$, $\beta$, $t$ are set to 39.9, 10, and 2 respectively, and $Q_{1}$ and $Q_{2}$ are randomly generated according to the Dirichlet distribution[38]. The experiments are performed on an NVIDIA GeForce RTX A6000 server GPU with CUDA 11.8.
Firstly, we compare our proposed contract-based incentive mechanism, which operates under information asymmetry, with other methods: a contract-based mechanism with complete information, a greedy scheme, and a random scheme. As illustrated in Fig. 4, we can find that our proposed contract scheme consistently outperforms the greedy and random schemes. However, for identical parameter settings, the contract-based mechanism with complete information yields higher performance than our model. This result highlights the disadvantage of information asymmetry, as the MLLM service provider gains fewer benefits without precise knowledge of healthcare data holders’ types. Although a complete information scenario allows the MLLM service provider to offer the optimal contract item to each data holder by knowing their exact type, it is not a realistic environment. In practice, even with complete information, a rational healthcare data holder may provide misleading information to manipulate rewards, ultimately reducing the subjective utility for the MLLM service provider. Thus, our proposed contract model, which handles asymmetric information, proves to be more reliable and practical, achieving the highest utility in real-world scenarios.
@@
In Fig. 5, we compare the performance of the GDM and DRL with Proximal Policy Optimization (DRL-PPO) in optimal contract design. Both models are capable of continuously acquiring rewards in complex and variable environments until convergence. Notably, the final test rewards of GDMs are significantly higher than those for DRL-PPO under identical parameter settings, allowing the MLLM service provider to consistently secure greater utilities. This is attributed to the fine-grained policy adjustments during the diffusion process, which effectively reduces the impact of randomness and noise[38]. Additionally, exploration through diffusion enhances the flexibility and robustness of the contract design policy, preventing it from falling into suboptimal solutions. Consequently, this superior performance demonstrates GDM’s ability to capture intricate patterns and connections among environmental observations, and it can effectively reduce the complexity of the relationship between healthcare data holders and the MLLM service provider.
@@
Figure 6: Optimal contracts designed by the GDM and DRL-PPO.
@@ VII Conclusion
In this paper, we studied the service quality issues of MLLMs and the design of incentive mechanisms for healthcare data management. We proposed a hybrid RAG-empowered medical MLLM framework based on cross-chain technologies to enhance data management in the healthcare industry. Specifically, we utilized a cross-chain structure comprising a main chain and multiple subchains to ensure the security of healthcare data. Additionally, we applied hybrid RAG with multi-modal information similarity metrics to retrieve similar healthcare data, thereby improving the quality of MLLM services. Then, we applied AoI to indirectly quantify the quality of healthcare data and utilized contract theory to incentivize healthcare data holders to contribute high-quality healthcare data with small AoI, thus enhancing the quality of MLLM services. Furthermore, we employed GDMs to generate the optimal contracts for efficient data sharing. Finally, numerical results demonstrated the effectiveness and reliability of our proposed framework and incentive mechanism. For future work, we aim to further improve the performance of our framework by incorporating more characteristics of multi-modal healthcare data and design a multi-dimensional contract model to address more complex healthcare industry environments.

### v2

@@ VI Numerical Results
In this section, we present extensive experiments to evaluate the performance of the proposed hybrid RAG-empowered MLLM framework for healthcare analysis, as well as the effectiveness of the proposed incentive mechanism. MLLM inference is conducted using Python 3.10.14 on an Intel Xeon(R) Gold 6133 CPU and an NVIDIA RTX A6000 GPU. For the implementation of GDM-based DRL algorithms, the primary parameter settings are detailed in Table III, with experiments run on an NVIDIA GeForce RTX A6000 server GPU using CUDA 11.8.
@@ VI-A Case Study of Hybrid RAG-empowered Medical MLLMs
We simulate a prototype of the hybrid RAG-empowered MMLMs with the support of LLaVA-Med[32] and llamaindex [^1] https://docs.llamaindex.ai/. As illustrated in Fig. 2, we present two examples to demonstrate the functionality and application of our framework. Upon receiving a task query with multi-modal healthcare data, the framework first retrieves the corresponding data from the respective modal database and performs a preliminary screening of $K$ results based on the cosine similarity between vectors. Next, hybrid RAG further refines these results using the MIS metric to identify the best matches, which are then used as inputs to MLLMs. Finally, the MLLMs process all this information to provide diagnostic outputs and personalized services according to the query.
We apply the criteria of Responsive Artificial Intelligence (RAI) [^2] https://www.microsoft.com/en-us/ai/responsible-ai to assess whether the outputs of MLLMs present potential risks related to morality, bias, and ethics. Additionally, we also assess the relationship between the output of MLLMs and task query with the Semantic Similarity (SS)[43], which reflects the diagnosis quality and serves as a crucial indicator for measuring the output of MLLMs. Due to the lack of evaluation benchmarks, we integrate LLM evaluators[44] with prompt engineering techniques [31] to measure the quality of data analysis for MLLMs under the method of GPT4-o, LLaVA, and LLaVA with hybrid RAG. The scoring is normalized to a range of $[0,1]$. Higher scores denote greater reliability and lack of bias in the MLLM output, along with a strong correlation to task query information. In contrast, lower scores signify a substantial gap from the anticipated results. Finally, we combine the RAI evaluation and SS into a unified score, known as the relative LLM score $\zeta$[45], which is calculated using the formula:
$\zeta=\lambda\cdot RAI+\nu\cdot SS,$ (20)
where $\lambda$ and $\nu$ are the weighting factors for RAI and SS, respectively. In our approach, we assign equal weights by setting $\lambda=0.5$ and $\nu=0.5$.
As shown in Fig. 3, we present the performance comparison between the proposed framework under different healthcare data cases. Our findings indicate that hybrid RAG enables LLaVA-Med to consistently score above $0.9$, particularly in X-ray cases from Users 1 and 2 with known etiologies, maintaining high-quality answers and stability. In contrast, other MLLMs exhibit reduced output quality due to the interference of disease factors. In the scenarios involving Users 3 and 4, who are normal without specific causes, MLLMs achieve high scores and deliver reasonable judgments. However, in the case of User 5, who is normal but has an X-ray that can easily be misjudged by a doctor, other MLLMs exhibit a higher misjudgment rate. In contrast, hybrid RAG continues to produce high-quality outputs by matching similar disease conditions. These cases illustrate that the retrieved data retrieved by hybrid RAG provides valuable information for answering questions. We summarize all scores in Table IV, which clearly demonstrates that hybrid RAG helps LLAVA-Med maintain consistently high scores, showcasing its strong performance across different scenarios. This indicates that hybrid RAG effectively considers the quality of retrieved information by utilizing the features of multi-modal data, including images and texts. The retrieved relevant healthcare data can aid MLLMs through contextual relationships, allowing MLLMs to deliver reliable and robust outputs owing to their powerful contextual learning capabilities.
@@ VI-B Performance of GDM-based Contract Theory Approach
Figure 6: Optimal contracts designed by the GDM and DRL-PPO.
In the proposed contract model, we employ an on-policy GDM algorithm within a double actor-critic framework for optimal contract design, and the specific settings of training hyperparameters are shown in Table III. In our setup, we consider $10$ healthcare data holders divided into two types, with $M=10$ and $K=2$. For the two types of healthcare data holders $\theta_{1}$ and $\theta_{2}$, values are randomly sampled from the intervals $[1,6]$ and $[13,18]$, respectively. Additionally, the maximum tolerance of AoI $\overline{A}_{max}$ is sampled randomly within the range of $[30,60]$. For the utility of the MLLM service provider, the parameters $\alpha$, $\beta$, $t$ are set to $39.9$, $10$, and $2$, respectively, and $Q_{1}$ and $Q_{2}$ are randomly generated according to the Dirichlet distribution[46].
Firstly, we compare our proposed contract-based incentive mechanism, which operates under information asymmetry, with other methods: a contract-based mechanism with complete information, a greedy scheme, and a random scheme. As illustrated in Fig. 4, we can find that our proposed contract scheme consistently outperforms the greedy and random schemes. However, for identical parameter settings, the contract-based mechanism with complete information yields higher performance than our model. This result highlights the disadvantage of information asymmetry, as the MLLM service provider gains fewer benefits without precise knowledge of the types of healthcare data holders. Although a complete information scenario allows the MLLM service provider to offer the optimal contract items to healthcare data holders by knowing their exact types, it is not a realistic environment. In practice, even with complete information, a rational healthcare data holder may provide misleading information to manipulate rewards, ultimately reducing the subjective utility for the MLLM service provider. Thus, our proposed contract model, which handles asymmetric information, proves to be more reliable and practical, achieving the highest utility in real-world scenarios.
In Fig. 5, we compare the performance of the GDM and DRL with Proximal Policy Optimization (DRL-PPO) in optimal contract design. Both models are capable of continuously acquiring rewards in complex and variable environments until convergence. Notably, the final test rewards of GDMs are significantly higher than those for DRL-PPO under identical parameter settings, allowing the MLLM service provider to consistently secure greater utilities. This is attributed to the fine-grained policy adjustments during the diffusion process, which effectively reduces the impact of randomness and noise[46]. Additionally, exploration through diffusion enhances the flexibility and robustness of the contract design policy, preventing it from falling into suboptimal solutions. Consequently, this superior performance demonstrates the ability of GDMs to capture intricate patterns and connections among environmental observations, and it can effectively reduce the complexity of the relationship between healthcare data holders and the MLLM service provider.
@@
VI-C Secure Block Verification Performance Analysis
To assess the security of the blockchain system, we evaluate the reputation value of hospitals. Each hospital’s associated subchain generates a block and broadcasts it to the relay chain for validation. If validated, the relay chain submits the block to the main chain linked to the health center. The health center then rewards each hospital according to their actions, applying a reputation-based bonus and penalty system.
As illustrated in Fig. 7, we use the Practical ByzantineFault Tolerance (PBFT) consensus algorithm to assess the security performance of the blockchain system. We consider that the subchains operate reliably and model the relay chain’s security performance as a random sampling problem with two potential outcomes, i.e., malicious delegates and well-behaved delegates[47]. When the number of malicious delegates is no greater than $(N-1)/3$, where $N$ represents the total number of delegates, the block verification process remains accurate[47]. Therefore, the probability of secure consensus, denoted as $P_{\mathrm{safety}}=\sum_{z=0}^{\lfloor N/3\rfloor}\binom{N}{z}p_{m}^{z}(1-p_{m})^{N-z}$, depends on $p_{m}$ that represents the probability of a delegate being malicious. Figure 7 shows that as the size of the relay chain increases, the security probability also rises, regardless of the likelihood of malicious delegates. This improvement is due to the larger number of well-behaved delegates involved in block validation, which strengthens security in the consensus process. Thus, the proposed blockchain system with the PBFT consensus algorithm supports secure and reliable data sharing by ensuring robust block verification.
Figure 7: Security probability at varying malicious miner probabilities.
@@ VII Conclusion
In this paper, we have studied the service quality issues of MLLMs and the design of incentive mechanisms for healthcare data management. We have proposed a hybrid RAG-empowered medical MLLM framework based on cross-chain technologies to enhance healthcare data management in IoMT. Specifically, we have utilized a cross-chain structure comprising a main chain and multiple subchains to ensure the security of healthcare data. Additionally, we have applied hybrid RAG with multi-modal information similarity metrics to retrieve similar healthcare data, thereby improving the quality of MLLM services. Then, we have applied AoI to indirectly quantify the quality of healthcare data and utilized contract theory to incentivize healthcare data holders to contribute high-quality healthcare data with small AoI, thus enhancing the quality of MLLM services. Furthermore, we have employed GDMs to generate the optimal contracts for efficient data sharing. Finally, numerical results demonstrate the effectiveness and reliability of our proposed framework and incentive mechanism. For future work, we aim to enhance our framework’s performance by integrating additional characteristics of multi-modal healthcare data and developing a multi-dimensional contract model to better address the complexities of IoMT environments.